# Online Makespan Scheduling under Scenarios

### v1

@@ 6 Conclusion and Outlook
Throughout, we have established several competitiveness results for Online Makespan Scheduling under Scenarios. The first main takeaway from our work is a competitiveness gap between two and at least three scenarios when we consider $m=2$ machines. This result draws a parallel to the known tractability results of several other problems, in which the line between easy and hard was drawn between two and three scenarios. Surely, the fact that the subsets of $\{1,2,3\}$ are not laminar, has been a deciding factor in our simple non-competitiveness result, as has been in several of the well-known results. Interestingly, restricting to the proxy competitive ratio as well as fixing the method of assigning double-scenario jobs essentially reduced the existence problem to the description of a polyhedron. We believe that similar techniques might prove useful for other cases and related problems as well.
@@
In light of our results, we may raise several open questions: In addition to the obvious task of finding tight bounds for all possibilities $(m,K)$ of the numbers of machines and scenarios, exploring the competitiveness as $m$ approaches to infinity in the weighted processing time case remains interesting. However, these problems have not been resolved even for $K=1$, whence it is fair to assume challenges in further progress. Another interesting research direction is to introduce migration, as seen in e.g. [migration]. In this model, we are allowed to revoke a bounded size of decisions in hindsight. Similarly, instead of minimizing the worst-case scenario, one could look into minimizing the average scenario or maximum regret.

### v2

@@ 7 Conclusion and Outlook
Throughout, we have established several competitiveness results for Online Makespan Scheduling under Scenarios. The first main takeaway from our work is a competitiveness gap between two and at least three scenarios when we consider $m=2$ machines. This result draws a parallel to the known tractability results of several other problems, in which the line between easy and hard was drawn between two and three scenarios. Surely, the fact that the subsets of $\{1,2,3\}$ are not laminar, has been a deciding factor in our simple non-competitiveness result, as has been in several of the well-known results. It is worth mentioning that restricting to the proxy competitive ratio as well as fixing the method of assigning double-scenario jobs essentially reduced the existence problem to the description of a polyhedron. We believe that this technique, if understood better, might prove useful for other cases and related problems as well.
@@
In light of our results, we may raise several open questions: In addition to the obvious task of finding tight bounds for all possibilities $(m,K)$ of the numbers of machines and scenarios, exploring the competitiveness as $m$ approaches to infinity in the weighted processing time case remains interesting. However, these problems have not been resolved even for $K=1$, whence it is fair to assume challenges in further progress. Another interesting research direction is to introduce migration, as seen in e.g. [migration]. In this model, we are allowed to revoke a bounded number of decisions in hindsight. Similarly, instead of minimizing the worst-case scenario, one could look into minimizing the average scenario or maximum regret.